OpenStudy (anonymous):

A 20% acid solution is mixed with a 70% acid solution to get 50 liters of a 40% solution. How much of the 70% solution is used?

OpenStudy (jhannybean):

@nincompoop

OpenStudy (anonymous):

x+y=50 .2x + .4y=50*.7 you would solve this like any system x is the amount in liters of the 20% solution and y is the amount of 40% solution

OpenStudy (anonymous):

do you know how to solve a linear system?

OpenStudy (anonymous):

i mixed that up the 2nd equation is .2x +.7y = 50 * .4

OpenStudy (anonymous):

x + y = 50 .2x +.7y = 50 * .4

OpenStudy (mathmath333):

this uses the concept of alligation and mixture  \(\Large \dfrac{a}{b}=\dfrac{70-40}{40-20}\\ \Large \dfrac{a}{b}=\dfrac{3}{2}\\ \Large \dfrac{a+b}{b}=\dfrac{3+2}{2}\\ \Large \dfrac{50}{b}=\dfrac{5}{2}\\\ \Large b=20 ~~litres\)

OpenStudy (anonymous):

thx